can you please do 30, 31, 37, 39b and show work?

30) A machine has a probability 0.03 of producing a defective glf b When 3 golf balls are manufactured by this machine, find the probability that none of the golf balls is defective 31) When three people are selected randomly, find the probability that their birthdays are all different 32) The mean credit card debt for college students in the U.S. is $2100. Is 2100 a parameter or statistic? Justify your ans wer 33) How many different outcomes are possible when 3 dice and 2 coins are tossed? 34) Determ ine whether the following events are a) disjoint b) independent or dependent A) Randomly selecting a male driver who weighs less than 140 pounds B) Randomly selecting a male driver who has more than three speeding tickets 35) (2 points each) Fl in the blank a) If the distribution of a random variable is skewed to the right, the value of the mean is than the value of the m edian s am ple is the one in which respondents them selves decide whether to be included in the sample and often reflects opinions of those with strong interests c) The running time of a random ly selected m ovie is an example of d) A researcher divides the population into three groups by their incomes: low, medium, and high variable Then he selects a random sample from each group. He used method in this study sam pling e) There are possible outcomes when a coin is tossed 6 times 36) a) A correlation between two variables does not im ply one is a_____of the other b) n a s tudy, data are collected from the past by going back in time sam pling method, every nth item is selected is a numerical measurement of a population e) A value is unusual if its z.score is more than or less than-____. f To apply the 68-95-99.7 principle, the distribution ofthe variable mut be g) When the effects of one factor cannot be separated from the effects of some other factors, the effects are said to be h) When the researcher controls the assignment of members to different groups, the study is study i) Truefalse: If the distribution is s kewed, we are usually better off with the mean than the m edian 37) A recent study finds that 34% of adult American males are obese. Bas ed on this study, wh at is the probability that a) all 3 are obese b) at least one is obese when 3 adult Am erican males are chosen random ly? 38) How m any different social security numbers with the last four digits 1234 are possible? 39) A random sam ple of 300 Mt. SAC students majoring in m ath or psychologyis selected. These students were asked whether they are happy with their m ajors. The results are summarized in the following table Happy Unhappy 100 Math 20 Ps ychology 10 170 a) When one student is selected randomly from this group, find the probability this student is happy with the choice of m ajor given that the student is a math major b) When one student is selected randomly from this group, find the probability this student is happy with the choice of major or the student is a math major c) When two students are chosen random ly, find the probability that they are both math m ajors 40) (3 points) Choose the data set with a sm aller standard deviation. Justify your answer
Answer #1

30)

The probability that a ball is not defective is

P(non defective) = 1 - P(defective) = 1 - 0.03 = 0.97

The probability that none of golf balls out of three is defective is

P(none of the golf ball is defective) = 0.97 *0.97 *0.97 = 0.912673

31)

Total number of ways in which three people can select their birthday is 365*365*365.

Number of ways three people can select different birthday is 365 *364 *363.

The probability that all three have different birthday is

P(all have 3 different birthday) = (365*364*363) / (365*365*365) = 0.9918